OpenStudy (anonymous):

Eliminate the parameter. x = t2 + 2, y = t2 - 4

OpenStudy (anonymous):

@phii

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

y = x - 6, x ≥ 1 y = x + 6, x ≥ 1 y = x2 - 6, x ≥ 1 y = x2 + 6, x ≥ 1

OpenStudy (anonymous):

\[\begin{cases}x=t^2+2\\y=t^2-4\end{cases}\] Since both \(x\) and \(y\) contain \(t^2\), I suggest solving one for \(t^2\), then substituting that into the other equation.

OpenStudy (anonymous):

\[x=t^2+2~\Rightarrow~t^2=x-2\] So, you have \[y=(x-2)-4\]

OpenStudy (anonymous):

y = x - 6

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

so its a? you rock